# Localizing Events in Videos with Multimodal Queries

### v1

@@ A.2.2 Prompt Engineering
1 | I have a caption {INPUT DATA}, adjust the {MODIFICATION TYPE} from {MODIFIED DETAIL} to {ORIGINAL DETAIL}. The revised caption should remain coherent and logical without introducing any additional details. | I have a caption {INPUT DATA}. Modify it by adding {NEW TYPE} {NEW DETAIL}. The revised caption should remain coherent and logical without introducing any other additional details. |
@@
2 | Read this {INPUT DATA}! Change the {MODIFICATION TYPE} from {MODIFIED DETAIL} to {ORIGINAL DETAIL}. Then, write a new caption that fits and doesn’t add new stuff. Only give the caption, no extra words. | Read this {INPUT DATA}! Add the {NEW TYPE} {NEW DETAIL} to it. Then, write a new caption that fits and doesn’t add new stuff. Only give the caption, no extra words. |
@@
3 | Here’s a caption {INPUT DATA}. Can you change {MODIFICATION TYPE} from {MODIFIED DETAIL} to {ORIGINAL DETAIL}? After that, make a new caption that makes sense and doesn’t add anything extra. Just write the caption, no explanations needed. | Here’s a caption {INPUT DATA}. Can you add {NEW TYPE} {NEW DETAIL}? After that, make a new caption that makes sense and doesn’t add anything extra. Just write the caption, no explanations needed. |
A.3 Extended Results
A.3.1 Model Performance with Different Prompts
We demonstrate the results of 3 prompts across different models and various metrics, including R1@0.5, R1@0.7, mAP@0.5, mAP@0.7, and Avg, showing both the Cap and Sum methods.
In Fig. 7, the results indicate that the performance is consistent across different metrics, demonstrating the robustness of the models when using Cap. The models generally maintain similar performance levels regardless of the specific metric, suggesting their stability and reliability in Cap. In contrast, Fig. 8 illustrates the performance of the same models but using Sum. It is evident that prompt 1 consistently outperforms prompts 2 and 3 across all metrics. This indicates that the models are more sensitive to the formulation of the prompt in Sum.
Method | original | ||||
@@ A.3.1 Model Performance with Different Prompts
A.3.2 Captioning Without Refinement Text V.S. Visual Query Encoding
We compare the model performance between Cap without refinement text and VisEnc, as shown in Tab. A.3.1. Both methods only use reference images as queries without refinement texts. Overall, Cap without refinement texts still significantly outperforms pure VisEnc, highlighting the effectiveness of image captioning. Additionally, TR-DETR and UVCOM perform best across all styles.
A.3.3 Original vs. Corrupted Text in Our Annotation
We have evaluated the model performance based on the original queries in QVHighlights and our corrupted texts to assess the significance of the refinement texts and the sensitivity of different models to natural language queries. [45] points out that the impact of natural language query may be minimal for some existing models, such as Moment-DETR. As shown in Tab. 8, Moment-DETR exhibits relatively smaller drops across all metrics, supporting this claim. On the other hand, the latest models, such as CG-DETR and TR-DETR, experience larger performance drops, indicating a higher sensitivity to query modifications. Furthermore, SeViLA is extremely sensitive to query modifications, which is shown by severe declines in performance across all evaluated metrics. Overall, the considerable performance decline across various models demonstrates that our modifications significantly affect the original queries. This also shows that our introduced refinement texts are not semantically trivial for localizing with multimodal queries.

### v3

@@ B.3 Prompt Engineering
| 1 | I have a caption {INPUT DATA}, adjust the {MODIFICATION TYPE} from {MODIFIED DETAIL} to {ORIGINAL DETAIL}. The revised caption should remain coherent and logical without introducing any additional details. | I have a caption {INPUT DATA}. Modify it by adding {NEW TYPE} {NEW DETAIL}. The revised caption should remain coherent and logical without introducing additional details. |
@@
| 2 | Read this {INPUT DATA}! Change the {MODIFICATION TYPE} from {MODIFIED DETAIL} to {ORIGINAL DETAIL}. Then, write a new caption that fits and doesn’t add new stuff. Only give the caption, no extra words. | Read this {INPUT DATA}! Add the {NEW TYPE} {NEW DETAIL} to it. Then, write a new caption that fits and doesn’t add new stuff. Only give the caption, no extra words. |
@@
| 3 | Here’s a caption {INPUT DATA}. Can you change {MODIFICATION TYPE} from {MODIFIED DETAIL} to {ORIGINAL DETAIL}? After that, make a new caption that makes sense and doesn’t add anything extra. Just write the caption; no explanations are needed. | Here’s a caption {INPUT DATA}. Can you add {NEW TYPE} {NEW DETAIL}? After that, make a new caption that makes sense and doesn’t add anything extra. Just write the caption, no explanations needed. |
Appendix C Extended Results
Due to the page limits, we appended additional experiments and analyses in this section.
C.1 Main Results for Other Metrics
We present the model performance in mAP in Tab. C.1 as an extension to Table 4.1. We find that the table aligns with the results stated in Sec. 5. Our SUIT strategy demonstrates good transferability to ICQ-Highlight. We highlight this in Fig. 10 on scribble images and show the performance gain with MQ-Sum(+SUIT) method.
{NiceTabular}llcccccccc[colortbl-like]
[…]
9-11
1
[…]
scribble
cartoon
cinematic
realistic
mAP@0.5 Avg. mAP@0.5 Avg. mAP@0.5 Avg. mAP@0.5 Avg.
VQ-Enc
Moment-DETR (2021) 14.95 6.67 16.51 7.21 17.00 7.39 17.41 7.66
QD-DETR (2023) 19.48 10.11 19.57 10.18 18.07 9.54 18.88 9.94
QD-DETR (2023) 18.22 9.74 14.31 7.30 15.18 7.45 14.71 7.66
EaTR (2023) 25.27 13.98 25.95 14.21 26.83 14.70 26.65 14.49
CG-DETR (2023) 30.24 15.57 30.78 15.70 30.07 15.48 30.98 15.83
TR-DETR (2024) 21.09 11.67 20.87 11.71 19.62 11.02 19.72 10.76
UMT (2022) 5.57 2.81 4.66 1.96 5.60 2.46 4.59 2.23
UniVTG (2023) 24.30 13.02 20.80 11.56 19.85 10.99 19.42 10.95
UVCOM (2023) 20.13 11.15 20.19 11.96 20.67 12.37 20.73 12.03
MQ-Cap
TimeChat (2024) 9.08 (± 0.6) 4.45 (± 0.4) 11.01 (± 0.9) 5.13 (± 0.5) 10.58 (± 0.7) 4.82 (± 1.0) 10.69 (± 1.0) 4.78 (± 0.2)
VTimeLLM (2024) 18.48 (± 1.0) 8.15 (± 0.5) 21.90 (± 0.3) 9.16 (± 0.1) 24.03 (±
0.5) 10.15 (± 0.3) 23.45 (± 0.7) 10.10 (± 0.1)
MQ-Sum
TimeChat (2024) 9.12 (± 0.4) 4.07 (± 0.2) 9.63 (± 1.7) 4.64 (± 0.7) 10.18 (± 1.2) 4.94 (± 0.9) 9.46 (± 1.8) 4.16 (± 1.3)
VTimeLLM (2024) 19.40 (± 1.4) 8.54 (± 0.4) 21.59 (± 0.8) 8.98 (± 0.4) 22.74 (± 0.3) 9.44 (± 0.3) 23.2 (± 1.6) 9.65 (± 0.7)
MQ-Sum
+ SUIT
Moment-DETR (2021) 49.46 (± 0.6) 28.36 (± 0.47) 49.01 (± 0.3) 28.0 (± 0.2) 49.32 (± 0.5) 28.07 (± 0.3) 48.39 (± 0.4) 27.34 (± 0.2)
QD-DETR (2023) 55.82 (± 0.2) 35.19 (± 0.1) 54.12 (± 0.5) 33.94 (± 0.2) 55.05 (± 0.2) 34.59 (± 0.2) 54.62 (± 0.2) 34.45 (± 0.2)
QD-DETR (2023) 54.71 (± 0.5) 35.29 (± 0.3) 54.20 (± 0.1) 35.48 (± 0.2) 54.05 (± 0.17) 35.2 (± 0.4) 53.14 (± 0.6) 34.54 (± 0.2)
EaTR (2023) 55.2 (± 0.7) 35.86 (± 0.4) 52.88 (± 0.2) 34.18 (± 0.2) 54.07 (± 0.7) 34.66 (± 0.1) 52.68 (± 0.3) 33.92 (± 0.4)
CG-DETR (2023) 55.6 (± 0.6) 36.16 (± 0.2) 55.5 (± 0.4) 35.47 (± 0.3) 55.93 (± 0.7) 35.85 (± 0.3) 55.34 (± 0.6) 35.43 (± 0.3)
TR-DETR (2024) 56.75 (± 0.4) 37.25 (± 0.2) 55.76 (± 0.2) 36.31 (± 0.1) 56.36 (± 0.5) 36.84 (± 0.5) 56.18 (± 0.3) 37.05 (± 0.3)
UMT (2022) 46.55 (± 0.3) 30.45 (± 0.3) 46.44 (± 0.6) 30.71 (± 0.3) 46.86 (± 0.4) 30.9 (± 0.3) 46.54 (± 0.2) 29.94 (± 0.2)
UniVTG (2023) 43.36 (± 0.4) 26.87 (± 0.2) 42.2 (± 0.4) 26.42 (± 0.2) 43.23 (± 0.5) 26.81 (± 0.3) 42.89 (± 0.58) 26.45 (± 0.4)
UVCOM (2023) 54.18 (± 0.3) 36.92 (± 0.4) 54.56 (± 0.3) 36.91 (± 0.1) 54.43 (± 0.4) 37.29 (± 0.1) 53.31 (± 0.5) 36.53 (± 0.2)
C.2 Model Performance on Different Refinement Text Types
We calculate the model performance on different subsets of refinement texts shown in Fig. 9. We conclude even though models have close performance across reference image styles, they show varied performance on different refinement text types across styles. For scribble style, models generally perform for “relation” better than other styles. For cartoon style, models demonstrate a more balanced performance across all types. The performance is notably higher for “environment” and “attribute” in cinematic style. Finally, for realistic style, the models yield better performance in “object” and “environment”.
| Method | original NLQ (Performance on QVHighlights) | ||||
@@ C.2 Model Performance on Different Refinement Text Types
C.3 MQ-based vs. NLQ-based Performance
We compare model performance on the MQ-based ICQ-Highlight and the original NLQ-based QVHighlight (results taken from the original papers) using Spearman’s rank correlation coefficient [69] on R1@0.5. For scribble, Spearman’s rank correlation coefficients are 0.89(MQ-Cap) and 0.93(MQ-Sum). The cartoon style yields coefficients of 0.98(MQ-Cap) and 0.94(MQ-Sum). The cinematic style shows coefficients of 0.93 for both MQ-Cap and MQ-Sum. Lastly, realistic has coefficients of 0.96(MQ-Cap) and 0.95(MQ-Sum). The high correlation scores indicate a strong positive correlation across benchmarks, suggesting queries of both benchmarks share the common semantics and yield the reliability of our benchmark.
C.4 Captioning Without Refinement Text vs. Visual Query Encoding
We compare the model performance between MQ-Cap without the revision step with refinement texts and VQ-Enc, as shown in Tab. C.6. Both methods only use reference images as queries without refinement texts. Overall, MQ-Cap without refinement texts still significantly outperforms pure VQ-Enc, highlighting the effectiveness of image captioning. Additionally, TR-DETR and UVCOM perform best across all styles.
C.5 Original NLQs (in QVHighlights) vs. Forged NLQs in ICQ-Highlight
We have evaluated the model performance based on the original NLQs in QVHighlights and our refinement texts introduced in MQs to assess the significance of the refinement texts and the sensitivity of different models to natural language queries. [60] points out that the impact of the NLQs may be minimal for some existing models, such as Moment-DETR. As shown in Tab. 8, Moment-DETR exhibits relatively smaller drops across all metrics, supporting this claim. On the other hand, the latest models, such as CG-DETR and TR-DETR, experience more significant performance drops, indicating a higher sensitivity to query modifications. Furthermore, SeViLA is extremely sensitive to query modifications, shown by severe performance declines across all evaluated metrics. Overall, the considerable performance decline across various models demonstrates that our modifications significantly affect the original queries. This also shows that our introduced refinement texts are not semantically trivial for localizing with multimodal queries.
C.6 Case Study: the Impact of Potential Generation Artifact
Along with the controlled experiment shown in Sec. 5.3, we conduct a qualitative case study with samples in the subsets and . We notice that generation artifacts usually do not change the image semantics and thus do not influence the caption dramatically, as shown in Fig. 11.
While collecting this subset, we noticed that AI-generated images become more prevalent on the Internet. This indicates that our generated dataset has a more realistic application and reflects the practical scenarios when users aim to locate events with generated images online. In addition, we find that generation artifacts do not pose significant issues in scribble and cartoon styles since the images are already simple.
@@ C.6 Case Study: the Impact of Potential Generation Artifact
| Type | Prompt | |||
| Object |
| Attributes |
| Actions |
| Environment |
| Relations |